a cone has a volume of 600 cubic inches and a radius of 6 inches. Find the height of the cone
Answer by oscargut(2103)

V = r^2(pi)h/3
600 = h(36(pi)/3)
12(pi)h = 600
(pi)h = 50
h = 50/pi
h = 15.915
If you need more help with similar questions, homework, assignments, you can contact me now at:
mthman@gmail.com
Thanks